OpenStudy (vijeya3):

A uniform chain of length 2 m is kept on the table such that a length of 60 cm hangs freely from the edge of the table.The total mass of the chain is 4 kg.What is the work done in pulling the entire chain?

OpenStudy (shootingstar28):

l and m are length and mass of part of chain hanging down. M and L are the total length and mass of the chain. M = 4 kg L = 2 m l = 0.6 m m = l / L * 4 = (0.6 / 2) * 4 = 1.2 kg The tension in the chain hanging down = mg = 12 N (g=10 m/s^2) The Force required = Tension The Displacement = 0.6 m W= Force * Displacement = 12 * 0.6 = 7.2 J The answer would be 7.2 J.

OpenStudy (irishboy123):

pulling it where? question incomplete, i reckon? and is there friction?

OpenStudy (vijeya3):

@ShootingStar28 I don't understand this part-m = l / L * 4 = (0.6 / 2) * 4 = 1.2 kg

OpenStudy (shootingstar28):

That is the chain.

OpenStudy (vijeya3):

Yeah..okay I don't understand the formula that u have applied here...I am not familiar with such a formula @ShootingStar28